Need all parts some hints are giving ! If you don’t know don’t answer need experts only
Question # 3 (A) The probability distribution function of a continuous variable is as follows: p(x) k x (1+x2) 0 G1Sx3+1) else where where k is a constant. (i) (ii) (iii) Find the value of, k Find the expected mean and standard deviation Determine the cumulative probability between (-1 that is, P(x+0.5)? x + 0.5); Answer-hints: (iii)i P(x 0.5) 15/16] Indicated below is a data set of eight numbers depicting random occurrences of an epoch x. The assigned numbers (1. 2,. 8) for the RV, x represent respectively, the directions (N, S, E, W, NE, NW, SE and SW) of a hurricane-track; and, the corresponding occurrence probabilities of x are: p(x) as listed. Data set B-I (B-1) 4 p(x)0.10 0.20 0.18 0.15 0.05 0.05 0.12 0.15 (i) Determine the expected mean of x (ii) Find the standard deviation of the distribution of x (ii) Decide the most probable direction of the hurricane based on the central-tendency Performance of the underlying statistics. (vi) Plot: CDF of x
Answer #1

(A)(i)

For a valid pdf sum of all probabilities must be equal to 1. Now,

8k

So,

8h

8 0.375

(ii)

The expected value is

E(X)p(a)d0.375x(1 +2)d0.375+ 471

Now,

1 1 E(x?) = 0.2 / 2.2p(z)dz = 0.3752.2(1 + x*)dx = 0.375 +

The variance of X is

Var(X) = E(X2)-[E(X)|-= 0.2-0 = 0.2

(iii)

P(Xleq +0.5)=int_{-1}^{0.5}p(x)dx=int_{-1}^{0.5}0.375(1+x^{2})dx=0.375left [ x+rac{x^{3}}{3} ight ]_{-1}^{0.5}=0.375left [ 0.5+rac{0.5^{3}}{3}+1+ rac{1^{3}}{3} ight ]=0.703125

(B-1)

Following table shows the calculations:

(i)

The expected value is :

E(X)=sum xP(X=x)=4.23

(ii)

The standard deviation is

sigma=sqrt{sum x^{2}P(X=x)-left [ sum xP(X=x) ight ]^{2}}=2.3574

(iii)

Since expected value is 4.23 so most probable direction is W.

(vi)

Following table shows the CDF:

13838351 ·46678 12855525 00110011 12345678

Following is the graph:

8 7 6 5 4 3 2 1 0 (x=>x)d